OpenStudy (anonymous):

Write an equation of the line that is parallel to the x-axis and whose y-intercept is : 5

OpenStudy (saifoo.khan):

Parallel to x axis means that it has the slope of Zero.

OpenStudy (saifoo.khan):

y = 5

OpenStudy (saifoo.khan):

y = 0(x) + 5 y = 0 + 5 y = 5

OpenStudy (anonymous):

oh okay lol , so its just the same answer?

OpenStudy (anonymous):

like if the question was 21 theen the answer would be y=21?

OpenStudy (saifoo.khan):

Yes, i elaborated.

OpenStudy (saifoo.khan):

yes.

OpenStudy (anonymous):

okayy :D

OpenStudy (saifoo.khan):

:D